# FAILURE OF LICHNEROWICZ-TYPE RESULT IN PARABOLIC GEOMETRIES OF REAL RANK AT LEAST 33

### v3

@@ 3 Lowest Weights
Suppose $\mathfrak{g}$ is noncomplex simple with real rank at least $3$ and fixed parabolic subalgebra. Let $\Omega\in H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ be a $\mathfrak{g}_{0}$ lowest weight, so that $\Omega\in V_{\beta,\gamma,\zeta}$ for some restricted roots $\beta,\gamma,\zeta.$ Then $\zeta\in-\hat{\Delta}^{+}$ and $\zeta\neq-\beta,-\gamma$.
@@
It follows from Theorem 3.6 that there is a lowest weight vector $\widetilde{\Omega}=(\eta_{\widetilde{\beta}})_{\flat}\wedge(\eta_{\widetilde{\gamma}})_{\flat}\otimes\eta_{\widetilde{\zeta}}$ such that $\beta,\gamma,\zeta$ are $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ restricted to $\mathfrak{a}.$ By Borel-Weil-Bott Theorem, we can assume without loss of generality that $\widetilde{\beta}=\alpha_{i},\widetilde{\gamma}=s_{i}(\alpha_{j}),$ and $\widetilde{\zeta}=-s_{i}s_{j}(\mu)$ for some $(ij)\in W^{\mathfrak{p}}(2),$ so $\beta=\alpha_{i}|_{\mathfrak{a}},$ $\gamma=s_{i}(\alpha_{j})|_{\mathfrak{a}},$ and $\zeta=-s_{i}s_{j}(\mu)|_{\mathfrak{a}}.$ It follows from Proposition 3.8(a) that $\gamma\in-\hat{\Delta}^{+}.$
It remains to show that
@@
By Theorem 3.10, $\Omega\in V_{\beta,\gamma,\zeta}$ for $\zeta\in-\hat{\Delta}^{+}$ and $\zeta\neq-\beta,-\gamma.$ The first condition implies $\mathrm{im}(\Omega)\subset\mathfrak{b}_{-}$. The second condition implies $\mathrm{im}(\Omega\wedge 1)\subset\ker\Omega.$ If $\zeta\in-\hat{\Delta}^{+}(\mathfrak{p}^{+})$ then $\mathrm{im}(\Omega)\subset\mathfrak{g}_{-}.$ On the other hand, if $\zeta\in\hat{\Delta}(\mathfrak{g}_{0})$ then $\mathrm{im}(\Omega)\subset\mathfrak{g}_{0}\cap\mathfrak{b}_{-}\subset\mathfrak{k}_{\Omega}$ because $\Omega$ is a $\mathfrak{g}_{0}$ lowest weight vector. Therefore $\mathrm{im}(\Omega)\subset\mathfrak{g}_{-}\oplus\mathfrak{k}_{\Omega}$, and $\Omega$ has the Kruglikov-The property. ∎
@@ 4.3 The case $(\mathfrak{sl}_{4}(\mathbb{H}),P_{2,6})$
For the case $(\mathfrak{sl}_{4}(\mathbb{H}),P_{2,6})$, we were unable to find an appropriate lowest weight vector and instead settled for a non-lowest weight harmonic seed whose Cartan geometry has a compact quotient admitting essential transformations. We will also write $P_{2,6}$ to represent the corresponding subalgebra of $\mathfrak{sl}_{8}.$ Let $v\in H^{2}_{\mathbb{C}}((\mathfrak{sl}_{8})_{-},\mathfrak{sl}_{8})_{+}$ be a $\mathfrak{g}_{0}$ lowest weight vector associated to $(21)\in W^{\mathfrak{p}}_{+}(2)$ by Borel-Weil-Bott theorem. Since
@@
$H_{\mathbb{C}}^{2}(\mathfrak{sl}_{8},P_{2,6})\cong H_{\mathbb{R}}(\mathfrak{sl}_{4}(\mathbb{H}),P_{2,6})^{\mathbb{C}},$
so that relative to the real form $\bigwedge^{2}(\mathfrak{sl}_{4}(\mathbb{H})_{-}^{*})\otimes\mathfrak{sl}_{4}(\mathbb{H})\leq\bigwedge(\mathfrak{sl}_{8})_{-}^{*}\otimes\mathfrak{sl}_{8},$ the real and imaginary parts of $\eta_{\alpha_{3}+\alpha_{4}}\cdot v$ are real harmonic elements. Either the real or imaginary part is nonzero, so assume without loss of generality that the real part $\Omega$ is nonzero. Let the real parts of the three terms of equation (7) be $\phi_{1},\phi_{2},\phi_{3}$, so $\Omega=\phi_{1}+\phi_{2}+\phi_{3}$. Then by Lemma 3.5, $\phi_{1}\in V_{\beta_{1}+\beta_{2},\beta_{1},-\beta_{2}-\beta_{3}}$ and $\phi_{2}\in V_{\beta_{1},\beta_{1}+\beta_{2},-\beta_{2}-\beta_{3}}$ and $\phi_{3}\in V_{\beta_{1},\beta_{1},-\beta_{3}}$ for $\beta_{1},\beta_{2},\beta_{3}$ the simple restricted roots of $\mathfrak{sl}_{4}(\mathbb{H})$. Then $\mathrm{im}(\Omega\wedge 1)\subset\ker\Omega$ and $\mathrm{im}(\Omega)\subset\mathfrak{g}_{-}$, so $\Omega$ has the Kruglikov-The property. Because $\mathrm{im}(\Omega)\leq\mathfrak{b_{-}}$, Theorem 2.12 implies $\Omega$ is a harmonic seed. The vector $\Omega$ is a restricted weight vector of weight $\tau:=2\beta_{1}-\beta_{3}.$ Then $\hat{I}=\{\beta_{1},\beta_{3}\}$ and $\I^\beta_{2}\in\hat{\Delta}^{0}\backslash\hat${I}, so $\beta_{2}^{\sharp}\in\mathfrak{g}_{0}^{ss}\cap\mathfrak{a}$. The restricted diagram has type $A_{3}.$ From
@@ 5 Main Theorem
We will start with an analysis of real Lie algebras admitting a complex structure. Let $\mathfrak{s}$ be a split real form, and let $\mathfrak{g}=(\mathfrak{s}^{\mathbb{C}})^{\mathbb{R}}$. The restricted roots of $\mathfrak{s}$ and $\mathfrak{g}$ are the same, and the restricted rootspaces of $\mathfrak{s}$ complexify to the restricted rootspaces of $\mathfrak{g}.$ By Proposition 2.8(a) there is a $\mathfrak{s}_{0}$ equivariant injection
@@
    Let $\mathfrak{k}_{\Omega^{\prime}}\leq\mathfrak{s}_{0}^{\mathbb{C}}$ be the stabilizer subalgebra of $\Omega^{\prime}.$ For $X\in\mathfrak{k}_{\Omega}$ and $U,V\in\mathfrak{s}_{-}$, we know that $\mathrm{ad}(X)(\Omega^{\prime})(U,V)=0.$ Since $\mathrm{ad}(X)(\Omega^{\prime})$ is $\mathbb{C}$-bilinear, $\mathrm{ad}(X)(\Omega^{\prime})=0$. Therefore $\mathfrak{k}_{\Omega}\leq\mathfrak{k}_{\Omega^{\prime}}$, and in fact $\mathfrak{k}_{\Omega}^{\mathbb{C}}\leq\mathfrak{k}_{\Omega^{\prime}}$. It follows that
    $\mathrm{im}(\Omega^{\prime})\leq(\mathfrak{s}_{-}\oplus\mathfrak{k}_{\Omega})^{\mathbb{C}}\leq\mathfrak{s}_{-}^{\mathbb{C}}\oplus\mathfrak{k}_{\Omega^{\prime}}.$
@@
By Corollary 2.10 there is an $\mathfrak{s}_{0}^{\mathbb{C}}$-equivariant injection
$H^{2}_{\mathbb{C}}(\mathfrak{s}_{-}^{\mathbb{C}},\mathfrak{s}^{\mathbb{C}})\hookrightarrow H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g}).$ (9)
For $\Omega^{\prime}\in H^{2}_{\mathbb{C}}(\mathfrak{s}_{-}^{\mathbb{C}},\mathfrak{s}^{\mathbb{C}}),$ let $\Omega^{\prime\prime}$ be the corresponding element in $H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$. Let $\mathfrak{b}^{\prime\prime}_{-}$ be the direct sum of the negative restricted rootspaces of $\mathfrak{g}.$
@@
    Let $\mathfrak{k}_{\Omega^{\prime\prime}}\leq\mathfrak{g}_{0}$ be the stabilizer subalgebra of $\Omega^{\prime\prime}.$ But $\mathfrak{g}_{0}=\mathfrak{s}_{0}^{\mathbb{C}}$ and $\mathfrak{k}_{\Omega^{\prime\prime}}=\mathfrak{k}_{\Omega^{\prime}}$. Then
    $\mathrm{im}(\Omega^{\prime\prime})=\mathrm{im}(\Omega^{\prime})\subset\mathfrak{s}_{-}^{\mathbb{C}}\oplus\mathfrak{k}_{\Omega^{\prime}}=\mathfrak{g}_{-}\oplus\mathfrak{k}_{\Omega^{\prime\prime}}.$
@@
    The negative restricted rootspaces of $\mathfrak{g}$ are equal to the negative rootspaces of $\mathfrak{s}^{\mathbb{C}}.$ Therefore
@@
If $\mathfrak{g}$ is complex, let $\mathfrak{s}$ be a split real form such that $\mathfrak{g}\cong(\mathfrak{s}^{\mathbb{C}})^{\mathbb{R}}.$ The parabolic subalgebra of $\mathfrak{g}$ is induced by a parabolic subalgebra of $\mathfrak{s}.$ Because $\mathfrak{s}$ is split, it cannot be isomorphic to $\mathfrak{sl}_{4}(\mathbb{H}).$ By Lemma 5.3, there exists a lowest $\mathfrak{g}_{0}$ weight vector $\Omega$ of $H^{2}_{\mathbb{R}}(\mathfrak{s}_{-},\mathfrak{s})$ of weight $\tau$ such that $\tau^{\sharp}$ is not scaling, and constants $a_{0}:=\alpha^{\sharp}+R$ and $c_{0}$ satisfying the required conditions. By Corollary 3.11, $\Omega$ satisfies the Kruglikov-The property and $\mathrm{im}(\Omega)\subset\mathfrak{b}_{-}.$ Applying the maps in equations (8) and (9) to $\Omega$ provides an element $\Omega^{\prime\prime}\in H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ of weight $\tau.$ By Proposition 5.1 and Proposition 5.2, $\Omega^{\prime\prime}$ satisfies the Kruglikov-The property and $\mathrm{im}(\Omega^{\prime\prime})\subset\mathfrak{b}^{\prime\prime}_{-}.$ Then by Theorem 2.12, $\Omega^{\prime\prime}$ is a harmonic seed.

### v4

@@ 4.2 Some Consequences
Suppose $\mathfrak{g}$ is noncomplex simple with real rank at least $3$ and fixed parabolic subalgebra. Let $\Omega\in H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ be a $\mathfrak{g}_{0}$ lowest weight, so that $\Omega\in\mathfrak{g}_{\beta,\gamma,\zeta}$ for some restricted roots $\beta,\gamma,\zeta.$ Then $\zeta\in-\hat{\Delta}^{+}$ and $\zeta\neq-\beta,-\gamma$.
@@
It follows from Theorem 4.7 that there is a lowest weight vector $\widetilde{\Omega}=(\eta_{\widetilde{\beta}})_{\flat}\wedge(\eta_{\widetilde{\gamma}})_{\flat}\otimes\eta_{\widetilde{\zeta}}$ such that $\beta,\gamma,\zeta$ are $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ restricted to $\mathfrak{a}.$ By Kostant’s Borel-Weil-Bott, we can assume without loss of generality that $\widetilde{\beta}=\alpha_{i},\widetilde{\gamma}=s_{i}(\alpha_{j}),$ and $\widetilde{\zeta}=-s_{i}s_{j}(\mu)$ for some $(ij)\in W^{\mathfrak{p}}(2),$ so $\beta=\alpha_{i}|_{\mathfrak{a}},$ $\gamma=s_{i}(\alpha_{j})|_{\mathfrak{a}},$ and $\zeta=-s_{i}s_{j}(\mu)|_{\mathfrak{a}}.$ It follows from Proposition 4.9(a) that $\gamma\in-\hat{\Delta}^{+}.$ It remains to show that
@@
By Theorem 4.11, $\Omega\in\mathfrak{g}_{\beta,\gamma,\zeta}$ for $\zeta\in-\hat{\Delta}^{+}$ and $\zeta\neq-\beta,-\gamma.$ The first condition implies $\mathrm{im}(\Omega)\subset\mathfrak{b}_{-}$. The second condition implies $\mathrm{im}(\Omega\wedge 1)\subset\ker\Omega.$ If $\zeta\in-\hat{\Delta}^{+}(\mathfrak{p}^{+})$ then $\mathrm{im}(\Omega)\subset\mathfrak{g}_{-}.$ On the other hand if $\zeta\in\hat{\Delta}(\mathfrak{g}_{0})$ then $\mathrm{im}(\Omega)\subset\mathfrak{g}_{0}\cap\mathfrak{b}_{-}$, and $\mathfrak{g}_{0}\cap\mathfrak{b}_{-}\subset\mathfrak{k}_{\Omega}$ because $\Omega$ is a $\mathfrak{g}_{0}$ lowest weight vector. Therefore $\mathrm{im}(\Omega)\subset\mathfrak{g}_{-}\oplus\mathfrak{k}_{\Omega}$, and $\Omega$ has the Kruglikov-The property. ∎
@@ 5.3 The case $(\mathfrak{sl}_{4}(\mathbb{H}),P_{2,6})$
For the case $(\mathfrak{sl}_{4}(\mathbb{H}),P_{2,6})$, we were unable to find an appropriate lowest weight vector and instead settled for a non-lowest weight harmonic seed whose Cartan geometry has a compact quotient admitting essential transformations. We will also write $P_{2,6}$ to represent the corresponding subalgebra of $\mathfrak{sl}_{8}.$ Let $v\in H^{2}_{\mathbb{C}}((\mathfrak{sl}_{8})_{-},\mathfrak{sl}_{8})_{+}$ be a $\mathfrak{g}_{0}$ lowest weight vector associated to $(21)\in W^{\mathfrak{p}}_{+}(2)$ by Kostant’s Borel-Weil-Bott. Since
@@
$H_{\mathbb{C}}^{2}((\mathfrak{sl}_{8})_{-},P_{2,6})\cong H_{\mathbb{R}}(\mathfrak{sl}_{4}(\mathbb{H})_{-},P_{2,6})^{\mathbb{C}},$
so that relative to the real form $\bigwedge^{2}(\mathfrak{sl}_{4}(\mathbb{H})_{-}^{*})\otimes\mathfrak{sl}_{4}(\mathbb{H})\leq\bigwedge(\mathfrak{sl}_{8})_{-}^{*}\otimes\mathfrak{sl}_{8},$ the real and imaginary parts of $\eta_{\alpha_{3}+\alpha_{4}}\cdot v$ are real harmonic elements. Either the real or imaginary part is nonzero, so assume without loss of generality that the real part $\Omega$ is nonzero. Let the real parts of the three terms of equation (3) be $\phi_{1},\phi_{2},\phi_{3}$, so $\Omega=\phi_{1}+\phi_{2}+\phi_{3}$. Then by Lemma 4.6, $\phi_{1}\in\mathfrak{sl}_{4}(\mathbb{H})_{\beta_{1}+\beta_{2},\beta_{1},-\beta_{2}-\beta_{3}}$ and $\phi_{2}\in\mathfrak{sl}_{4}(\mathbb{H})_{\beta_{1},\beta_{1}+\beta_{2},-\beta_{2}-\beta_{3}}$ and $\phi_{3}\in\mathfrak{sl}_{4}(\mathbb{H})_{\beta_{1},\beta_{1},-\beta_{3}}$ for $\beta_{1},\beta_{2},\beta_{3}$ the simple restricted roots of $\mathfrak{sl}_{4}(\mathbb{H})$. Then $\mathrm{im}(\Omega\wedge 1)\subset\ker\Omega$ and $\mathrm{im}(\Omega)\subset\mathfrak{g}_{-}$, so $\Omega$ has the Kruglikov-The property. Because $\mathrm{im}(\Omega)\leq\mathfrak{b_{-}}$, Theorem 3.7 implies $\Omega$ is a harmonic seed. The vector $\Omega$ is a restricted weight vector of weight $\tau:=2\beta_{1}-\beta_{3}.$ Then $\hat{I}=\{\beta_{1},\beta_{3}\}$ and $\I^\beta_{2}\in\hat{\Delta}^{0}\backslash\hat${I}, so $\beta_{2}^{\sharp}\in\mathfrak{g}_{0}^{ss}\cap\mathfrak{a}$. The restricted diagram has type $A_{3}.$ From
@@ 6 Main Theorem
We will start with an analysis of real Lie algebras admitting a complex structure. Let $\mathfrak{s}$ be a split real form, and consider the underlying real Lie algebra of $\mathfrak{s}^{\mathbb{C}}$. The restricted roots of $\mathfrak{s}$ and $\mathfrak{s}^{\mathbb{C}}$ are the same, and the restricted rootspaces of $\mathfrak{s}$ complexify to the restricted rootspaces of $\mathfrak{s}^{\mathbb{C}}.$ By Proposition 3.4(b) there is a $\mathfrak{s}_{0}$-equivariant injection
@@
    Let $\mathfrak{k}_{\Omega^{\prime}}\leq\mathfrak{s}_{0}^{\mathbb{C}}$ be the stabilizer subalgebra of $\Omega^{\prime}.$ For $X\in\mathfrak{k}_{\Omega}$ and $U,V\in\mathfrak{s}_{-}$, we know that
    $\mathrm{ad}(X)(\Omega^{\prime})(U,V)=\mathrm{ad}(X)(\Omega)(U,V)=0.$
    Since $\mathrm{ad}(X)(\Omega^{\prime})$ is $\mathbb{C}$-bilinear, $\mathrm{ad}(X)(\Omega^{\prime})=0$. Therefore $\mathfrak{k}_{\Omega}\leq\mathfrak{k}_{\Omega^{\prime}}$, and in fact $\mathfrak{k}_{\Omega}^{\mathbb{C}}\leq\mathfrak{k}_{\Omega^{\prime}}$. It follows that
    $\mathrm{im}(\Omega^{\prime})=\mathrm{im}(\Omega)^{\mathbb{C}}\leq(\mathfrak{s}_{-}\oplus\mathfrak{k}_{\Omega})^{\mathbb{C}}\leq\mathfrak{s}_{-}^{\mathbb{C}}\oplus\mathfrak{k}_{\Omega^{\prime}}.$
@@
By Proposition 3.5 there is an $\mathfrak{s}_{0}^{\mathbb{C}}$-equivariant injection
$H^{2}_{\mathbb{C}}(\mathfrak{s}_{-}^{\mathbb{C}},\mathfrak{s}^{\mathbb{C}})\hookrightarrow H^{2}_{\mathbb{R}}(\mathfrak{s}_{-}^{\mathbb{C}},\mathfrak{s}^{\mathbb{C}}).$ (5)
For $\Omega^{\prime}\in H^{2}_{\mathbb{C}}(\mathfrak{s}_{-}^{\mathbb{C}},\mathfrak{s}^{\mathbb{C}}),$ let $\Omega^{\prime\prime}$ be the corresponding element in $H^{2}_{\mathbb{R}}(\mathfrak{s}^{\mathbb{C}}_{-},\mathfrak{s}^{\mathbb{C}})$. Let $\mathfrak{b}^{\prime\prime}_{-}$ be the direct sum of the negative restricted rootspaces of $\mathfrak{s}^{\mathbb{C}}.$
@@
    Let $\mathfrak{k}_{\Omega^{\prime\prime}}\leq\mathfrak{s}_{0}^{\mathbb{C}}$ be the stabilizer subalgebra of $\Omega^{\prime\prime}.$ But $\mathfrak{k}_{\Omega^{\prime\prime}}=\mathfrak{k}_{\Omega^{\prime}}$. Then
    $\mathrm{im}(\Omega^{\prime\prime})=\mathrm{im}(\Omega^{\prime})\subset\mathfrak{s}_{-}^{\mathbb{C}}\oplus\mathfrak{k}_{\Omega^{\prime}}=\mathfrak{s}_{-}^{\mathbb{C}}\oplus\mathfrak{k}_{\Omega^{\prime\prime}}.$
@@
    The negative restricted rootspaces of $\mathfrak{s}^{\mathbb{C}}$ are equal to the negative rootspaces of $\mathfrak{s}^{\mathbb{C}}.$ Therefore
@@
If $\mathfrak{g}$ is the underlying real Lie algebra of a complex Lie algebra, let $\mathfrak{s}$ be a split real form such that $\mathfrak{g}\cong\mathfrak{s}^{\mathbb{C}}.$ Any parabolic subalgebra of $\mathfrak{g}$ is induced by a parabolic subalgebra of $\mathfrak{s}.$ Because $\mathfrak{s}$ is split, it cannot be isomorphic to $\mathfrak{sl}_{4}(\mathbb{H}).$ By Lemma 6.3, there exists a lowest $\mathfrak{g}_{0}$ weight vector $\Omega$ of $H^{2}_{\mathbb{R}}(\mathfrak{s}_{-},\mathfrak{s})$ of weight $\tau$ such that $\tau^{\sharp}$ is not scaling, and constants $a_{0}:=\alpha^{\sharp}+R$ and $c_{0}$ satisfying the required conditions. By Corollary 4.12, $\Omega$ satisfies the Kruglikov-The property and $\mathrm{im}(\Omega)\subset\mathfrak{b}_{-}.$ Applying the maps in equations (4) and (5) to $\Omega$ provides an element $\Omega^{\prime\prime}\in H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ of weight $\tau.$ By Proposition 6.1 and Proposition 6.2, $\Omega^{\prime\prime}$ satisfies the Kruglikov-The property and $\mathrm{im}(\Omega^{\prime\prime})\subset\mathfrak{b}^{\prime\prime}_{-}.$ Then by Theorem 3.7, $\Omega^{\prime\prime}$ is a harmonic seed.